# T2VEval: T2V-generated Videos Benchmark Dataset and Objective Evaluation Method

### v3

@@ Abstract
Recent advances in text-to-video (T2V) technology, as demonstrated by models such as Runway Gen-3, Pika, Sora, and Kling, have significantly broadened the applicability and popularity of the technology. This progress has created a growing demand for accurate quality assessment metrics to evaluate the perceptual quality of T2V-generated videos and optimize video generation models. However, assessing the quality of text-to-video outputs remain challenging due to the presence of highly complex distortions, such as unnatural actions and phenomena that defy human cognition. To address these challenges, we constructed T2VEval-Bench, a multi-dimensional benchmark dataset for text-to-video quality evaluation, comprising 148 textual prompts and 1,783 videos generated by 13 T2V models. To ensure a comprehensive evaluation, we scored each video on four dimensions in the subjective experiment, which are overall impression, text-video consistency, realness, and technical quality. Based on T2VEval-Bench, we developed T2VEval, a multi-branch fusion scheme for T2V quality evaluation. T2VEval assesses videos across three branches: text-video consistency, realness, and technical quality. Using an attention-based fusion module, T2VEval effectively integrates features from each branch and predicts scores with the aid of a large language model. Additionally, we implemented a progressive training strategy, enabling each branch to learn targeted knowledge while maintaining synergy with the others. Experimental results demonstrate that T2VEval achieves state-of-the-art performance across multiple metrics.
@@ I INTRODUCTION
Based on the T2VEval-Bench dataset, we further designed T2VEval, an objective evaluation method for T2V-generated. Previous works, such as FETV [liu2024fetv], EvalCrafter [liu2024evalcrafter], and VBench [huang2024vbench], have developed dedicated evaluation metrics for each evaluation dimension. However, these metrics generally lack broad applicability, and they directly employ models trained on datasets with human-made videos, thereby overlooking the inherent distributional differences between human-made videos and AI-generated videos. Other studies like T2VQA[kou2024subjective] and AIGC-VQA [lu2024aigc] have implemented end-to-end evaluations of videos, but they only consider video quality and text-video alignment. Therefore, we designed a multi-branch fusion evaluation scheme, which includes a technical quality encoding branch, a video realness encoding branch, and a text-video alignment encoding branch. The technical quality encoder is implemented based on the Swin-3D network [liu2022video], the video realness encoder utilizes the ConvLexNet-3D network [liu2022convnet], and the text-video alignment is achieved using BLIP [li2022blip]. Considering that technical quality and realness rely solely on visual perception, we prioritize the fusion of quality and realness features before integrating them with text-video alignment features. Additionally, we adopted a progressive training strategy that ensures the independent effectiveness of each branch while promoting collaboration among the branches.
Our contributions are summarized as follows:
  • We constructed a multi-dimensional benchmark dataset named T2VEval-Bench. The construction of this dataset involves four key aspects: the division of evaluation dimensions, the compilation of textual prompts, video generation, and the collection of subjective scores. We define four scoring dimensions (overall impression, text-video consistency, realness, and technical quality) to reflect common flaws in text-to-video generation. Furthermore, we collected a large number of videos generated by T2V models for evaluation. Finally, we conducted subjective experiments to gather over 60,000 subjective scores, and perform a series of data analyses. T2VEval-Bench provided important data support for subsequent objective evaluation algorithms.
  • We introduced a multi-branch fusion framework for text-to-video quality evaluation called T2VEval, which incorporates a technical quality encoding branch, a video realness encoding branch, and a text-video alignment encoding branch. In terms of model architecture, a progressive training strategy was adopted: we first fused visually pertinent features (technical quality and realness) before integrating the text-video alignment branch. This design ensures the independent effectiveness of each branch while maintaining overall synergy, leading to more precise and objective assessments of defects in T2V-generated videos.
  • Our designed T2V objective evaluation metric achieved state-of-the-art performance on both T2VEval-Bench and T2VQA-DB[kou2024subjective]. Experimental results show that T2VEval effectively assesses the performance of different T2V models and has broad potential applications.
@@ II-A Subjective evaluation and dataset for T2V Generation
All of the above works focus on constructing general subjective evaluation benchmarks for text-to-video. At the same time, there are also works that conduct subjective evaluations for specific video content types or evaluation angles. Chen et al. [chen2024gaia] constructed the GAIA dataset for action quality assessment of generated videos from the perspective of actions in the generated videos, covering various full-body, hand, and facial actions. TC-Bench [feng2024tc] focuses on evaluating the temporal composition and consistency of each element in the generated videos. T2VBench [ji2024t2vbench] and DEVIL [liao2024evaluation] evaluate from the perspective of the dynamic degree of text-to-video. T2V-CompBench [sun2024t2v] evaluates T2V models from seven dimensions such as dynamic attribute binding, spatial relationship, and generation computing ability in terms of the ability of T2V models to combine different objects, attributes, and actions. In addition, Miao et al. [miao2024t2vsafetybench] paid special attention to the security risks of text-to-video. They pointed out that the generated videos may contain illegal or unethical content, and previous work lacked a quantitative understanding of security. Therefore, the authors introduced T2VSafetyBench to evaluate T2V models in 12 aspects such as violence, gore, discrimination, political sensitivity, copyright, and trademark infringement.
Figure 3: The Prompt generation strategy used in our work.
@@ III-B Compilation and Screening of Prompts
Figure 4: Prompt content category distribution and example of prompt-video pairs.
TABLE I: Detailed information about the models included in this article, including frame rate, duration, resolution, abilities and release date of the version used in the experiment. *Sora has not yet opened its interface, and only 48 published videos participated in our evaluation.
Model & Version Duration FPS Resolution Abilities
Sora* 8-60s 30fps $1280\times 720$ or $1920\times 1080$ T2V
Runway Gen-3 Alpha 8s 24fps $2816\times 1536$ T2V & I2V & V2V
Kling 1.0 5s 30fps $1280\times 720$ T2V & I2V
Dreamina 1.0 12s 8fps $1680\times 720$ T2V & I2V
TypeMovie 32s 24fps $1024\times 576$ T2V & I2V
Pixeling 5s 24fps $1024\times 576$ T2V & I2V
Dream Machine 1.0 5s 24fps $1360\times 752$ T2V & I2V
Vega AI 4s 12fps $1024\times 576$ T2V & I2V
PixVerse V2 4s 18fps $1408\times 768$ or $1024\times 576$ T2V & I2V
Pika 1.0 3s 24fps $1280\times 720$ T2V & I2V
Show-1 3.63s 8fps $576\times 320$ T2V
VideoCrafter 1.6s 10fps $512\times 320$ T2V & I2V
Open-Sora 1.2 4.25s 24fps $1280\times 720$ T2V & I2V & V2V
@@
Figure 5: Distribution of subjective scores in each dimension.
TABLE II: The final ranking of the participating models. The red font represents the highest score, the blue font represents the second highest score, and the green font represents the lowest score. *Sora’s scores only includes 48 prompts’ MOSs, which may have certain limitations.
@@ III-C Video Generation
We select 13 most representative and advanced T2V models, including Sora, Pika 1.0, Runway Gen-3 Alpha, Open-Sora 1.2, PixVerse V2, Show-1, VideoCrafter, Dreamina 1.0, Kling 1.0, Dream Machine 1.0, Pixeling, and TypeMovie. Details of these models are provided in Table I. We generated a total of 1,783 videos including 48 publicly available videos from Sora and 1,735 videos generated independently. The distribution of video quantities across each category is illustrated in Figure 4(a).
@@ III-D Collection of Subjective Scores
Figure 6: The radar maps of each model under different prompt’s categories. Considering that Sora only had 48 demo videos, which may affect the reliability of evaluation, we did not include it in our analysis.
@@
After removing outliers, we normalized the data and calculate the Mean Opinion Score (MOS) for each video. For a video generated by model $i$ on prompt $j$ in dimension $k$, denoted as $MOS_{ijk}$. The final score for each video was obtained by averaging its scores across all evaluators:
[…]
where $N$is the number of evaluators’ scores after excluding outliers. Finally, we obtained a multi-dimensional benchmark dataset containing 1783 videos, each with MOS scores in four dimensions.
Figure 5 shows the MOS distribution for each dimension, demonstrating that the scores cover nearly the entire possible range and follow a normal distribution. This confirms the comprehensiveness and validity of the subjective scores.
III-E Subjective Evaluation Results
We comprehensively analyzed the subjective evaluation results and computed the T2V model’s scores across four evaluation dimensions to summarize the performance of each model. Table II provides the scores of the 13 models across all evaluation dimensions, while Figure 7 visualizes the score distribution using a violin plot. Notably, Sora achieves the highest scores in all dimensions, exhibiting a highly concentrated and stable distribution. However, it is important to note that Sora’s ranking is based solely on its 48 demo videos, which introduces certain limitations to the evaluation.
Figure 7: The Violin distribution maps of the model’s subjective score in each evaluation dimension
@@ III-E Subjective Evaluation Results
Finally, we analyzed the performance of these models in each category, and the results are shown in Figure 6. Runway Gen-3 consistently outperforms other models across all prompt categories, while Open-Sora 1.2, VideoCrafter, and Show-1 exhibit the weakest performance. This aligns with the conclusions presented in Table II. Furthermore, it is evident that regardless of the prompt category, the scores for the realness dimension are generally lower than those of other dimensions. This highlights that the primary challenge faced by current T2V models lies in accurately understanding and representing the objective laws of the real world, including physical principles and cognitive commonsense knowledge. Specifically, these models struggle to accurately understand and represent the physiological structures (e.g., facial features, limbs), movement patterns (e.g., running, dancing), and physical laws (e.g., sliding, dripping) of organisms in the real world, as illustrated in Figure 8.
@@ IV-D Integrated Scoring prediction Module
TABLE III: Performance comparison of different models. Red: the best
@@ V-B Performance Comparison
Table III compares the performance of T2VEval with other methods. The results demonstrate that our method achieves state-of-the-art performance on T2VEval-Bench, surpassing the second-best method by nearly 4%. The performance of zero-shot models is notably poor, which may be attributed to their sole focus on text-video alignment and inability to capture temporal information between video frames. Furthermore, these methods are typically trained and tested on datasets with human-made videos, overlooking the domain gap between AI-generated videos and human-made videos. Although VQA models achieve higher scores, they still fall significantly short of T2VQA and T2VEval, further underscoring the necessity of multi-branch fusion evaluation framework evaluation frameworks.
@@
Additionally, we fine-tuned and tested T2VEval on the T2VQA-DB dataset, following the same data partitioning method as T2VEval-Bench. The results, averaged across 10 partitions, are also presented in Table III. These experiments confirm that T2VEval also achieves state-of-the-art performance in this dataset.
@@ V-C Ablation Studies
To systematically evaluate the contribution of core architectural components in T2VEval, we conducted comprehensive ablation studies. Specifically, in Table IV, we conducted experiments to evaluate (i) different vision fidelity fusion mechanisms and (ii) differential pre-training strategies applied to the realness encoder. Additionally, in Table V, we explored the contribution and effect of each branch, aiming to gain a deeper understanding of how these components interact and influence the overall performance of T2VEval.
@@ V-C1 Ablation Experiment of Vision Fidelity Fusion Strategy
T2VEval employs a cross-attention mechanism to fuse technical quality and realness features. In previous studies, alternative fusion methods, such as feature summation and concatenation, have also been utilized. To evaluate their effectiveness, we compared the cross-attention fusion strategy against addition and concatenation methods (Addition is denoted as T2VEval-Add, and concatenation is denoted as T2VEval-Cat). As shown in Table IV, the cross-attention method significantly outperforms the other two approaches.
@@
Results are shown in Table IV. These results may be due to the fact that realness and technical quality are completely dependent on visual content and have a stronger task-dependence. In contrast, due to the existence of the text modality, fusing visual-related features (whether realness feature or technical quality feature) with consistency features first may lead to interference among various modal features.
TABLE IV: Performance comparison of ablation studies. Red: the best
@@
TABLE V: Ablation study on the usage of technical quality branch, realness branch and alignment branch in T2VEval. Red: the best
@@ V-C2 Ablation Experiment of Realness Branch Pre-train Strategy
When designing the realness evaluation branch, we introduced the pre-training task of AI-generated video detection, instead of simply using ImageNet pre-trained weights or random initialization. We replaced the pre-trained weights of Conv3D with ImageNet weights (denoted as T2VEval-ImageNet) and random initialization (denoted as T2VEval-Random) respectively, and fine-tuned them in the same training environment. The experimental results are shown in Table IV. In addition, we further used Grad-CAM to analyze the impact of different pre-training strategies on the model’s focus area, as shown in Figure 11. The figure demonstrates the activation maps of the 3D network loaded with different pre-trained weights.
@@ V-C3 Ablation Experiment of The Effectiveness of Each Branch and Training Strategy
In this section, we performed ablation experiments to evaluate the effectiveness of each branch and the divide-and-conquer training strategy employed. Specifically, we removed each of the three branches of the model and conducted fine-tuning tests on the T2VEval-Bench dataset. The dataset partitioning method remains consistent with the previous experiments. The results, presented in Table V, demonstrate that removing any branch leads to varying degrees of performance degradation, confirming the importance and effectiveness of the three dimensions assessed in the T2V quality evaluation task. Notably, removing the technical quality branch has the smallest impact on model performance, suggesting that when evaluating text-to-video generation systems, focusing solely on technical quality may prove insufficient. A comprehensive assessment framework should additionally prioritize ”content quality evaluation”, encompassing two critical dimensions we proposed: (1) visual realness and (2) text-video alignment.
Additionally, we evaluated the impact of the divide-and-conquer training strategy we proposed by comparing it with a single-stage training approach where the entire model is trained at once using the overall impression MOS scores (denoted as T2VEval-Once). The results, shown in Table IV, indicate that our divide-and-conquer training strategy outperforms the single-stage approach. This finding highlights that the divide-and-conquer strategy effectively facilitates cooperation between branches, leading to improved overall performance.

### v5

# T2VEval: Benchmark Dataset and Objective Evaluation Method for T2V-generated Videos

@@ Abstract
Recent advances in text-to-video (T2V) technology, as demonstrated by models such as Runway Gen-3, Pika, Sora, and Kling, have significantly broadened the applicability and popularity of the technology. This progress has created a growing demand for accurate quality assessment metrics to evaluate the perceptual quality of T2V-generated videos and optimize video generation models. However, assessing the quality of text-to-video outputs remain challenging due to the presence of highly complex distortions, such as unnatural actions and phenomena that defy human cognition. To address these challenges, we constructed T2VEval-Bench, a multi-dimensional benchmark dataset for text-to-video quality evaluation, which contains 148 textual prompts and 1,783 videos generated by 13 T2V models. To ensure a comprehensive evaluation, we scored each video on four dimensions in the subjective experiment, which are overall impression, text-video consistency, realness, and technical quality. Based on T2VEval-Bench, we developed T2VEval, a multi-branch fusion scheme for T2V quality evaluation. T2VEval assesses videos across three branches: text-video consistency, realness, and technical quality. Using an attention-based fusion module, T2VEval effectively integrates features from each branch and predicts scores with the aid of a large language model. Additionally, we implemented a divide-and-conquer training strategy, enabling each branch to learn targeted knowledge while maintaining synergy with the others. Experimental results demonstrate that T2VEval achieves state-of-the-art performance across multiple metrics.
@@ I INTRODUCTION
Based on the T2VEval-Bench dataset, we further designed T2VEval, an objective evaluation method for T2V-generated. Previous works, such as FETV [liu2024fetv], EvalCrafter [liu2024evalcrafter], and VBench [huang2024vbench], have developed dedicated evaluation metrics for each evaluation dimension. However, these metrics generally lack broad applicability, and they directly employ models trained on datasets with human-made videos, thereby overlooking the inherent distributional differences between human-made videos and AI-generated videos. Other studies like T2VQA[kou2024subjective] and AIGC-VQA [lu2024aigc] have implemented end-to-end evaluations of videos, but they only consider video quality and text-video alignment. Therefore, we designed a multi-branch fusion evaluation scheme, which includes a technical quality encoding branch, a video realness encoding branch, and a text-video alignment encoding branch. The technical quality encoder is implemented based on the Swin-3D network [liu2022video], the video realness encoder utilizes the ConvLexNet-3D network [liu2022convnet], and the text-video alignment is achieved using the BLIP network [li2022blip]. Considering that technical quality and realness rely solely on visual perception, we prioritize the fusion of quality and realness features before integrating them with text-video alignment features. Additionally, we adopted a progressive training strategy that ensures the independent effectiveness of each branch while promoting collaboration among the branches. Experimental results show that T2VEval effectively assesses the performance of different T2V-generated videos and has broad potential applications, achieved state-of-the-art performance on both T2VEval-Bench and T2VQA-DB[kou2024subjective].
@@ II-A Subjective evaluation and dataset for T2V Generation
All of the above works focus on constructing general subjective evaluation benchmarks for text-to-video generation. At the same time, there are also works that conduct subjective evaluations for specific video content types or evaluation angles. Chen et al. [chen2024gaia] constructed the GAIA dataset for action quality assessment of generated videos from the perspective of actions in the generated videos, covering various full-body, hand, and facial actions. TC-Bench [feng2024tc] focuses on evaluating the temporal composition and consistency of each element in the generated videos. T2VBench [ji2024t2vbench] and DEVIL [liao2024evaluation] evaluate from the perspective of the dynamic degree of text-to-video. T2V-CompBench [sun2024t2v] evaluates T2V models from seven dimensions such as dynamic attribute binding, spatial relationship, and generation computing ability in terms of the ability of T2V models to combine different objects, attributes, and actions. In addition, Miao et al. [miao2024t2vsafetybench] paid special attention to the security risks of text-to-video. They pointed out that the generated videos may contain illegal or unethical content, and previous work lacked a quantitative understanding of security. Therefore, the authors introduced T2VSafetyBench to evaluate T2V models in 12 aspects such as violence, gore, discrimination, political sensitivity, copyright, and trademark infringement.
Figure 3: (a) The Prompt generation strategy used in our work. (b) The word cloud of the prompts used in T2VEval-Bench
@@ III-B Compilation and Screening of Prompts
Figure 4: Prompt content category distribution and example of prompt-video pairs. (a) Category-wise distribution pie chart of T2V-generated video counts in our T2VEval-Bench. (b) Sample entries from the T2VEval-Bench containing video frames and corresponding textual prompts.
@@
TABLE I: The rankings of the models being evaluated. The red font represents the highest score, the blue font represents the second highest score, and the green font represents the lowest score. *Sora’s scores only includes 48 prompts’ MOSs, which may have certain limitations.
@@ III-C Video Generation
We select 13 most representative and advanced T2V models, including Sora, Pika 1.0, Runway Gen-3 Alpha, Open-Sora 1.2, PixVerse V2, Show-1, VideoCrafter, Dreamina 1.0, Kling 1.0, Dream Machine 1.0, Pixeling, and TypeMovie. Details of these models are provided in Supplementary Table S1. We generated a total of 1,783 videos including 48 publicly available videos from Sora and 1,735 videos generated independently. The distribution of video quantities across each category is illustrated in Figure 4(a).
@@ III-D Collection of Subjective Scores
After removing outliers, we normalized the data and calculate the Mean Opinion Score (MOS) for each video. Let the video generated by model $i$ under prompt $j$ in dimension $k$ be $\textit{Video}_{ijk}$. If the score given by evaluator $l$ to $\textit{Video}_{ijk}$ is $S_{ijkl}$, then the Mean Opinion Score $\textit{MOS}_{ijk}$ of $\textit{Video}_{ijk}$ is:
$\textit{MOS}_{ijk}=\frac{1}{N}\sum_{l=1}^{N}S_{ijkl}$ (1)
where $N$ is the number of evaluators for $\textit{Video}_{ijk}$ after excluding outliers. Following the same method, the MOS score for each video in every dimension can be obtained.
Supplementary Figure S1 shows the MOS distribution for each dimension, demonstrating that the scores cover nearly the entire possible range and follow a normal distribution. This confirms the comprehensiveness and validity of the subjective scores.
III-E Analysis of T2V Model Performance
To analyze the performance of T2V models, we calculated the MOS scores of each model in every dimension. Let the mean opinion score of model $i$ in dimension $k$ be $\textit{MOS}_{ik}$, then the $\textit{MOS}_{ik}$ can be calculated as follows:
$\textit{MOS}_{ik}=\frac{1}{M}\sum_{j=1}^{M}\textit{MOS}_{ijk}$ (2)
where $M$ is the number of prompts. Table I provides the MOS scores of the 13 models across all evaluation dimensions, while Supplementary Figure S3 visualizes the score distribution using a violin plot. Notably, Sora achieves the highest scores in all dimensions, exhibiting a highly concentrated and stable distribution. However, it is important to note that Sora’s ranking is based solely on its 48 demo videos, which introduces certain limitations to the evaluation.
@@ III-E Analysis of T2V Model Performance
Finally, we analyzed the performance of these models in each category, and the results are shown in Supplementary Figure S2. Runway Gen-3 consistently outperforms other models across all prompt categories, while Open-Sora 1.2, VideoCrafter, and Show-1 exhibit the weakest performance. This aligns with the conclusions presented in Table I. Furthermore, it is evident that regardless of the prompt category, the scores for the realness dimension are generally lower than those of other dimensions. This highlights that the primary challenge faced by current T2V models lies in accurately understanding and representing the objective laws of the real world, including physical principles and cognitive commonsense knowledge. Specifically, these models struggle to accurately understand and represent the physiological structures (e.g., facial features, limbs), movement patterns (e.g., running, dancing), and physical laws (e.g., sliding, dripping) of organisms in the real world, as illustrated in Figure 5.
@@ IV-D Integrated Scoring prediction Module
TABLE II: Performance comparison of different models. Red: the best
@@ V-B Performance Comparison
Table II compares the performance of T2VEval with other methods. The results demonstrate that our method achieves state-of-the-art performance on T2VEval-Bench, surpassing the second-best method by nearly 4%. The performance of zero-shot models is notably poor, which may be attributed to their sole focus on text-video alignment and inability to capture temporal information between video frames. Furthermore, these methods are typically trained and tested on datasets with human-made videos, overlooking the domain gap between AI-generated videos and human-made videos. Although VQA models achieve higher scores, they still fall significantly short of T2VQA and T2VEval, further underscoring the necessity of multi-branch fusion evaluation framework evaluation frameworks.
@@
Additionally, we fine-tuned and tested T2VEval on the T2VQA-DB dataset, following the same data partitioning method as T2VEval-Bench. The results, averaged across 10 partitions, are also presented in Table II. These experiments confirm that T2VEval also achieves state-of-the-art performance in this dataset.
@@ V-C Ablation Studies
To systematically evaluate the contribution of core architectural components in T2VEval, we conducted comprehensive ablation studies. Specifically, in Table III, we conducted experiments to evaluate (i) different vision fidelity fusion mechanisms and (ii) differential pre-training strategies applied to the realness encoder. Additionally, in Table IV, we explored the contribution and effect of each branch, aiming to gain a deeper understanding of how these components interact and influence the overall performance of T2VEval.
@@ V-C1 Ablation Experiment of Vision Fidelity Fusion Strategy
T2VEval employs a cross-attention mechanism to fuse technical quality and realness features. In previous studies, alternative fusion methods, such as feature summation and concatenation, have also been utilized. To evaluate their effectiveness, we compared the cross-attention fusion strategy against addition and concatenation methods (Addition is denoted as T2VEval-Add, and concatenation is denoted as T2VEval-Cat). As shown in Table III, the cross-attention method significantly outperforms the other two approaches.
@@
Results are shown in Table III. These results may be due to the fact that realness and technical quality are completely dependent on visual content and have a stronger task-dependence. In contrast, due to the existence of the text modality, fusing visual-related features (whether realness feature or technical quality feature) with consistency features first may lead to interference among various modal features.
TABLE III: Performance comparison of ablation studies. Red: the best
@@
TABLE IV: Ablation study on the usage of technical quality branch, realness branch and alignment branch in T2VEval. Red: the best
@@ V-C2 Ablation Experiment of Realness Branch Pre-train Strategy
When designing the realness evaluation branch, we introduced the pre-training task of AI-generated video detection, instead of simply using ImageNet pre-trained weights or random initialization. We replaced the pre-trained weights of Conv3D with ImageNet weights (denoted as T2VEval-ImageNet) and random initialization (denoted as T2VEval-Random) respectively, and fine-tuned them in the same training environment. The experimental results are shown in Table III. In addition, we further used Grad-CAM to analyze the impact of different pre-training strategies on the model’s focus area, as shown in Figure 8. The figure demonstrates the activation maps of the 3D network loaded with different pre-trained weights.
@@ V-C3 Ablation Experiment of The Effectiveness of Each Branch and Training Strategy
In this section, we performed ablation experiments to evaluate the effectiveness of each branch and the divide-and-conquer training strategy employed. Specifically, we removed each of the three branches of the model and conducted fine-tuning tests on the T2VEval-Bench dataset. The dataset partitioning method remains consistent with the previous experiments. The results, presented in Table IV, demonstrate that removing any branch leads to varying degrees of performance degradation, confirming the importance and effectiveness of the three dimensions assessed in the T2V quality evaluation task. Notably, removing the technical quality branch has the smallest impact on model performance, suggesting that when evaluating text-to-video generation systems, focusing solely on technical quality may prove insufficient. A comprehensive assessment framework should additionally prioritize ”content quality evaluation”, encompassing two critical dimensions we proposed: (1) visual realness and (2) text-video alignment.
Additionally, we evaluated the impact of the divide-and-conquer training strategy we proposed by comparing it with a single-stage training approach where the entire model is trained at once using the overall impression MOS scores (denoted as T2VEval-Once). The results, shown in Table III, indicate that our divide-and-conquer training strategy outperforms the single-stage approach. This finding highlights that the divide-and-conquer strategy effectively facilitates cooperation between branches, leading to improved overall performance.